OpenStudy (anonymous):

Can someone tell me if I'm correct? VERY MUCH APPRECIATED (-1)(x^2) - 3/2x^69 - square root of 2.4 IS a polynomial. 1(x^2) - 3x^69/2 IS a polynomial. (-1)(x^2) - 3x^68 - square root of 2.4(x) IS a polynomial. 1(x^2) - 3x^69 - square root of x IS a polynomial. 1(x^2) - 3x^-96 IS NOT a polynomial.

OpenStudy (amistre64):

give your reasoning ....

OpenStudy (amistre64):

or, how would you define a polynomial?

OpenStudy (anonymous):

To be honest I was guessing. I know polynomials can't have a negative exponent or a variable as a denominator ?

OpenStudy (amistre64):

good, they also cant have anyting like square roots and such of a variable .... if memory serves, the coefficients are not restricted tho, or are they?

OpenStudy (amistre64):

and for "square root of", just type sqrt(x) its more readily understood

OpenStudy (amistre64):

(-1)(x^2) - 3/2x^69 - sqrt(2.4) ^ ^ is this under the fraction? 1(x^2) - 3x^69/2 ^ ^ good (-1)(x^2) - 3x^68 - square root of 2.4(x) is this sqrt(2.4x) ? or, x sqrt(2.4)? 1(x^2) - 3x^69 - sqrt(x) ^^ trouble 1(x^2) - 3x^-96 ^^ trouble

OpenStudy (amistre64):

hopefully that second one isnt x^(69/2)

OpenStudy (anonymous):

the ^69 is over the 3/2

OpenStudy (anonymous):

it's 2.4 times x

OpenStudy (amistre64):

\[\frac32x^{69} \text{ is a good term then}\]

OpenStudy (amistre64):

the the x is getting sqrted, then theres trouble

OpenStudy (amistre64):

*if the x is getting ..

OpenStudy (anonymous):

so the first on IS a polynomial?

OpenStudy (amistre64):

if ive read your post correctly, then the first one looks fine

OpenStudy (anonymous):

Here, can I take a screen shot of the question and attach it? would that be easier?

OpenStudy (amistre64):

that would be fine

OpenStudy (anonymous):

Ok give me a minute please and thank you :D

OpenStudy (amistre64):

much better :)

OpenStudy (anonymous):

I figured as much

OpenStudy (amistre64):

now, from the information ive already stated; can you tell me (by number) which ones you beleieve to be polys?

OpenStudy (anonymous):

1, 2, and 4 I believe

OpenStudy (amistre64):

spose by letter is fine .... dint see them there for some reason :)

OpenStudy (anonymous):

maybe 3

OpenStudy (amistre64):

does 2 have a fraction as an exponent on the x?

OpenStudy (anonymous):

Yes.

OpenStudy (amistre64):

does 69/2 simplify to an integer?

OpenStudy (anonymous):

So it isn't a poly

OpenStudy (anonymous):

69 is like 34. something ?

OpenStudy (amistre64):

then its not an integer result and is bad for poly definitions

OpenStudy (anonymous):

so b is NOT a poly

OpenStudy (amistre64):

b is NOT a poly

OpenStudy (amistre64):

... too many options for D, can a poly have a sqrt(x) in it?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

? I think lol

OpenStudy (amistre64):

\[\Large\sqrt x =x^{1/2}\] can we have exponents that are not integers?

OpenStudy (amistre64):

just remember that if x has anything other than an exponent and a coefficient acting on it, its bad news

OpenStudy (amistre64):

sin(x) is bad news ln(x) is bad news e^(x) is bad news sqrt(x) is bad news if the exponents is not a positive integer, ... its bad news as well

OpenStudy (amistre64):

from this, I can only see 2 options that are polynomials by definition

OpenStudy (amistre64):

\[(-1)\cdot x^2-\frac32x^{69}-\sqrt{2.4}~;~good \] \[1\cdot x^2-3\cancel {x^{69/2}}\] \[(-1)\cdot x^2-3 {x^{68}}-\sqrt{2.4}\cdot x~:~good\] \[1\cdot x^2-3 {x^{69}}-\cancel{\sqrt{x}}\] \[1\cdot x^2-3 \cancel{{x^{-69}}}\]

OpenStudy (anonymous):

So is it A and D?

OpenStudy (amistre64):

uh, no. its A and C ... barring any typos that is

OpenStudy (anonymous):

Ohh ok. Well thank you you've really helped a lot

OpenStudy (amistre64):

youre welcome, and good luck :)